OpenStudy (anonymous):

MEDALS + FAN!!!!!!!! Can ANY1 HELP ME WITH MY MIDDLE SCHOOL HOME WORK PLEASE JUST 2 OR 1 QUESTIONS????

OpenStudy (anonymous):

What are the questions?

OpenStudy (anonymous):

@abtster

OpenStudy (anonymous):

OK How many \(\frac{ 3 }{4 } \) bags can you make from a 5 pound bag 5 pounds = \[\frac{5 \text{ pound bag} }{3/4}=5 \times \frac{4}{3} \text{of} \frac{3}{4} \text{pound bags} = \frac{20}{3} = 6\frac{2}{3} \text{of} \frac{3}{4} \text{pound bags}\] So you can make 6 bags and have on only partially filled

OpenStudy (anonymous):

The second one: \[ \frac{ 3 }{ 4 } \times 8 =\frac{ 24 }{ 4 }=6\]

OpenStudy (anonymous):

if you need help doing work imagine this, 8 is the same as 8/1, so to make it easier do 3/4 * 8/1 and just multiply straight across, so find 3* 8, and 4 * 1, and you will get 24/4